OpenStudy (anonymous):

Solve -15x+10-3x=-6x+8-4x

OpenStudy (anonymous):

leave all the x's on one side and all the number values on the other side.

OpenStudy (yanasidlinskiy):

First step, for this equation is to combine like terms on both sides. You get: -18x+10= -10x+8 -------Now you need to get x by itself. Either x would work. +10x +10x -8x+10=8 -10 -10 -8x = -2 -----Divide by -8 x=4

OpenStudy (yanasidlinskiy):

Makes sense?

OpenStudy (anonymous):

thxs